# Phonon Induced Contrast in Matter Wave Interferometer

### v1

@@ VI Numerical Results
In this section, we numerically explore contrast reduction due to phonon wave packets from the interactions given in Eq. (7). Since we do not want the magnetic field gradient, which controls the spatial superposition state, to be too large ($\eta_{\text{b}}\leq 10^{6}$) T/m, see such gradients can be achieved in Machluf et al. (2013); Mamin et al. (2012), thus the magnetic field gradient determines the maximum spatial separation distances $\Delta X_{\text{m}}\sim 10^{-6}-10^{-3}$m for different masses. For simplicity, here we consider phonons as longitudinal elastic waves, $\omega_{q}=cq$ Jishi, Radi A (2013), and treat them as continuous with a minimum frequency (fundamental frequency) $\omega_{0}={\pi c}/{L}$. Besides, we take the eigenvector $Q_{q}^{i}=1$ (recall that in solid state physics, these vectors are sinusoidal functions) as an upper limit. We also define the moment when the diamond wavepacket reaches its maximum separation distance $\Delta X_{\text{m}}$ as time $t=0$. Thus, $t_{1}=-2\tau_{\text{a}}-\tau_{\text{f}}$, and $t_{6}=-t_{1}=2\tau_{\text{a}}+\tau_{\text{f}}$. To compute Eq. (25), we need to convert Eq. (10) and Eq. (11) into respective transfer functions. Here, we first provide their forms in the temporal domain as differences on both sides of the interferometer arms,
$\Delta F_{s}(t)$ $=\mu(S^{\text{R}}-S^{\text{L}})\eta_{\text{b}}=2\mu\eta_{\text{b}}$ (27)
$\Delta F_{\text{dia}}(t)$ $=N(F_{c}^{R}-F_{c}^{L})$
$=\frac{\chi_{\rho}M}{\mu_{0}}\eta_{b}^{2}\left(X_{c}^{\text{R}}-X_{c}^{\text{L}}\right).$
The differential force for spin-phonon interaction is easily understood in (27). These are the constant differences between the left and right interferometer arms at different time stages. Recall that $\eta_{\text{b}}$ is the maximum value of the magnetic field gradient. For the difference generated by diamagnetic repulsion, we have made an assumption: the diamagnetic repulsion experienced by the diamond as a bulk is averaged over each of its particles at its CoM, and we neglect the differences in diamagnetism between carbon atoms and nitrogen defects. Therefore, the summation over $i$ in $f_{q}$ of diamagnetic repulsion for each particle becomes the diamagnetic repulsion experienced by the central particle (averaged) multiplied by the number of particles $N$. Thereafter, we can obtain their corresponding transfer functions,
@@
In Fig. 5, we show the contrast reduction resulting from spin-magnetic field and diamagnetic interactions in a setup with $\Delta t=1$s. The horizontal axis represents the maximum separation distance of the diamond’s centre of mass in the spatial superposition state. The black and red lines compare diamonds of different masses, solid and long dash lines represent results at low temperature (4 K), while dotted and short dash lines correspond to room temperature (300 K). The results indicate that the contrast loss induced by the spin-magnetic field coupling is dominant, while the effect of diamagnetic interaction is significantly smaller. The source of these differences can be easily understood by comparisons in (3031). The contrast reduction caused by spin-magnetic field-phonon interaction is suppressed by $\omega_{q}^{3}$, whereas the contrast loss caused by phonon-diamagnetic repulsion is suppressed by $\omega_{q}^{7}$. 555 Since the main contribution of contrast loss comes from the fundamental frequency $\omega_{0}\sim c/L$. One can consider two limits here, one is $\hbar\omega_{0}\ll k_{\text{B}}T$ and the other is $\hbar\omega_{0}\gg k_{\text{B}}T$. For the former limit, the square of the transfer function will yield the maximum value of $|\Gamma(\omega_{0})|^{2}\sim 1$, and Eqs. (30) and (31) become, respectively, $-\text{Log}~{}C_{\text{spin}}\approx\frac{2k_{\text{B}}T}{\hbar^{2}}\frac{M\Delta X^{2}_{\text{m}}}{\tau_{\text{a}}^{4}\omega_{0}^{4}}\sim\frac{2k_{\text{B}}T}{\hbar^{2}}\rho\left(\frac{\Delta X_{\text{m}}}{c^{2}\tau_{a}^{2}}\right)^{2}L^{7},$ $-\text{Log}~{}C_{\text{dia}}\approx\frac{k_{\text{B}}T\chi_{\rho}^{2}}{8\hbar^{2}\mu_{0}^{2}\mu^{4}}\frac{M^{5}\Delta X_{\text{m}}^{6}}{\tau_{\text{a}}^{12}\omega_{0}^{8}}\sim\frac{k_{\text{B}}T\chi_{\rho}^{2}\rho^{5}}{8\mu_{0}^{2}\mu^{4}\hbar^{2}}\left(\frac{\Delta X_{\text{m}}^{3}}{c^{4}\tau_{\text{a}}^{6}}\right)^{2}L^{23}.$ In the other limit, when $\hbar\omega_{0}\gg k_{\text{B}}T$ the hyperbolic function goes $\coth\frac{1}{2}\frac{\hbar\omega_{0}}{k_{\text{B}}T}\rightarrow 1.$ The contrast reductions from the two interactions become as, $-\text{Log}~{}C_{\text{spin}}\approx\frac{M\Delta X^{2}_{\text{m}}}{\hbar\tau_{\text{a}}^{4}\omega_{0}^{3}}\sim\frac{\rho}{\hbar c^{3}}\left(\frac{\Delta X_{\text{m}}L^{3}}{\tau_{\text{a}}^{2}}\right)^{2},$ $-\text{Log}~{}C_{\text{dia}}\approx\frac{\chi_{\rho}^{2}}{16\mu_{0}^{2}\mu^{4}\hbar}\frac{M^{5}\Delta X^{6}_{\text{m}}}{\tau_{\text{a}}^{12}\omega_{0}^{7}}\sim\frac{\chi_{\rho}^{2}}{16\mu_{0}^{2}\mu^{4}\hbar}\frac{\rho^{5}}{c^{7}}\left(\frac{\Delta X^{3}_{\text{m}}L^{11}}{\tau_{\text{a}}^{6}}\right)^{2}.$ In the above two limits, we have used the relation: $M=\rho L^{3}$, where $\rho$ is the density of test masses and $L$ is the length. In either of the limits, one can always conclude that more rigid material (which means faster sound speed) and smaller size of the test masses will produce less reduction in contrast. The diamond-like crystal will be an example of an ideal crystal for the QGEM experiment.
@@
Moreover, room temperature does not drastically alter the outcomes for either of them. Additionally, it can be observed that larger diamond masses lead to more phonon contrast reduction, which is quite understandable. Larger mass implies a lower range of phonon frequencies. From Eqs. (25) and (28), it can be seen that these $\omega_{q}$ terms appear in the denominators. From Eq.(25), we can see that the acceleration protocol profoundly impacts the contrast of the phonon wave packets. In Fig. 6, we use a density plot to detail the contrast reduction resulting from different acceleration protocols. In the four subplots, we highlight the black curve, where the magnetic field gradient is precisely $10^{6}$ T/m, and in the region above it, the duration time is longer, thus requiring a smaller magnetic field gradient. In Fig. 6(a) and (b), we study the contrast of a diamond with a mass of $M=2.25\times 10^{-14}$ kg under different maximum separation distances $\Delta X_{\text{m}}$ and duration times. In the QGEM, the separation distance of the CoM wave packets should be sufficiently large (e.g., $\Delta x\sim 10^{-4}$ m). We can observe that under a one-second flight time, the contrast loss resulting from spin-magnetic field coupling and diamagnetic interaction ranges from $10^{-9}$ to $10^{-6}$ and $10^{-33}$ to $10^{-26}$, respectively. In (c) and (d), we set the CoM’s maximum separation distance to $10^{-4}$ m and explore the contrast loss for different diamond masses and duration times.
Consistent with intuition, larger masses lead to larger volumes, thus involving more phonon modes. A larger mass requires larger acceleration to create a large spatial superposition, hence a large force to be imparted on the diamond. The larger force leads to a significant loss, in contrast. For instance, at $\Delta t=0.05$ s, a diamond with a mass of $10^{-12}$ kg will experience a contrast reduction of approximately $1\%$ due to spin-magnetic field coupling. However, within the current experimental constraints where the SG magnetic field is constrained to be less than $10^{6}{\rm T/m}$ Machluf et al. (2013); Mamin et al. (2012), the contrast loss induced by both couplings, e.g. spin and induced diamagnetic terms, is negligible for QGEM.
@@ VII Discussion
In this paper, we have analyzed the contrast loss due to phonon in creating spatial superposition of a nanoscaled diamond embedded with a single NV centre based on a Stern-Gaelach interferometer. During the process of creating a spatial superposition state using the coupling between the spin and the linear magnetic field, the spin particle in the lattice chain generates a pair of forces (10) of opposite directions and equal magnitudes due to their own spin superposition state. Similar to the Humpty-Dumpty problem at the centre of mass of the diamond, the spin particles on both sides of the interferometer arm have different vibrational states in the lattice, resulting in differences in the quantum states of the phonons. This also leads to imperfections in the overlap of phonon wave packets. The other interaction arises from the diamagnetic properties of diamond. Each atom in the lattice chain produces a diamagnetic force (11) when subjected to an external magnetic field. This diamagnetic force is of the same direction but has a different magnitude on both sides of the interferometer arm. Similarly, due to this difference, the quantum states of phonons on both sides are also different, resulting in contrast reduction. In this paper, we use a linear magnetic field (8), which is widely used in atomic interferometers. Furthermore, this linearity greatly simplifies the calculation of the motion state of spin particles. That is, the splitting force is independent of trajectories and the equilibrium position. The situation is more complex regarding the interaction between diamagnetic force and phonons. Since the diamagnetic term in the Hamiltonian includes $B(X_{i})^{2}$, the force on the lattice is now coupled with its trajectory. This external field mediated coupling between phonon and CoM’s trajectory is common in the study of optical cooling of the motion state of diamond centre. In this work, we typically studied the latter situation and made comparison of contrast reduction between these two interactions. Specifically, we assumed that the phonons were initially at a specific temperature $T$, and we utilized the Wigner representation to express and compute the evolution of the quantum state of the phonons. Our results indicate that among these two sources contributing to the contrast loss in the phonon degrees of freedom, the effect of the spin-magnetic field coupling is dominant. The contrast loss induced by the diamagnetic term is smaller than the former, as shown in Fig. 5. This counter-intuitive result can be directly understood from (30) and (31). For the spin-magnetic field coupling, which does not depend on the trajectory, the contrast reduction is suppressed by the phonon frequency to the power of three. Meanwhile, the contrast loss due to diamagnetic interaction is suppressed by the phonon frequency to the power of seven. The latter case directly arises from its coupling with the trajectory, as the diamagnetic force is a function of time, and its variation is adiabatic. In the seminal work Henkel and Folman (2022, 2023), this is also referred to as the degree of adiabaticity. However, overall, the contrast reduction induced by both interactions is minimal. This is good news for matter-wave interferometers. There are two main reasons for this result. One is that we limit the gradient of the magnetic field. The most important reason, however, is that we are studying diamonds of very small size ($10^{-20}\sim 10^{-14}$ kg) compared to the speed of sound $c$ multiplied by the duration time (on the order of microseconds and 1 second for QGEM). Since the frequencies of the phonons are very high (tens of gigahertz to several terahertz), their contribution to contrast loss is tiny. However, if we prefer a different material where the speed of sound is much smaller than that of the diamond, we would see the loss of contrast occur. However, it also depends on how we create the superposition; hence, it depends on the details of the transfer function. Acknowledgements: SB thanks EPSRC grants EP/R029075/1, EP/X009467/1, and ST/W006227/1. Q.X and R.Z are supported by the China Scholarship Council (CSC).
Appendix A Initial thermal equilibrium mixed state

### v2

@@ VI Numerical Results
In this section, we numerically explore contrast reduction due to phonon wave packets from the interactions given in Eq. (8). Since we do not want the magnetic field gradient, which controls the spatial superposition size, to be too large (i.e., $\eta_{\text{b}}\leq 10^{6}$ T/m, see such gradients can be achieved in Machluf et al. (2013); Mamin et al. (2012)), thus the maximum spatial separation distances $\Delta X_{\text{m}}\sim 10^{-6}-10^{-3}$m for different masses. For simplicity, here we consider phonons as longitudinal elastic waves, $\omega_{q}=cq$ Jishi, Radi A (2013), and treat them as continuous with a minimum frequency (fundamental frequency) $\omega_{0}={\pi c}/{L}$. Besides, we take the eigenvector $Q_{q}^{i}=1$ (recall that in solid state physics, these vectors are sinusoidal functions) as an upper limit. We also define the moment when the diamond wavepacket reaches its maximum separation distance $\Delta X_{\text{m}}$ as time $t=0$. Thus, $t_{1}=-2\tau_{\text{a}}-\tau_{\text{f}}$, and $t_{6}=-t_{1}=2\tau_{\text{a}}+\tau_{\text{f}}$.
To compute Eq. (26), we need to convert Eq. (14) and Eq. (15) into respective transfer functions. Here, we first provide their forms in the temporal domain as differences on both sides of the interferometer arms,
$\Delta F_{s}(t)$ $=\mu(S^{\text{R}}-S^{\text{L}})\eta_{\text{b}}=2\mu\eta_{\text{b}},$ (28)
$\Delta F_{\text{dia}}(t)$ $=\sum_{i}\left(F_{i,\text{dia}}^{\text{R}}-F_{i,\text{dia}}^{\text{L}}\right)$
$=\frac{\chi_{\rho}M}{\mu_{0}}\eta_{b}^{2}\left(X_{c}^{\text{R}}-X_{c}^{\text{L}}\right)$
The differential force for spin-phonon interaction is easily understood in (28). These are the constant differences between the left and right interferometer arms at different time stages. Recall that $\eta_{\text{b}}$ is the maximum value of the magnetic field gradient.
Likewise, for the diamagnetism-phonon coupling, due to the diamond wave packet being in a spatial superposition, there is also a difference of the diamagnetic repulsion generated by the atoms located on the two sides of the interferometer arms (for example, the $i^{\text{th}}$ atom with positions $X_{i}^{\text{L,R}}$).
Thereafter, we can obtain their corresponding transfer functions,
@@
Figure 5: The figure shows the phonon-induced contrast loss caused by spin-phonon (solid lines) and diamagnetism-phonon (dash and dot lines) interactions in the expected scenario of the QGEM experiment (with a duration time of approximately one second). The horizontal axis $\Delta X_{\text{m}}$ is the maximum separation distance between diamond wave packets on either sides of the interferometer. From Eqs. (3132), the contrast loss caused by phonon modes will be zero when $\Delta X_{\text{m}}=0$ indicating no spatial separation. In the figure, the contrast loss from heavier mass $M=10^{-14}$kg (four black lines) is correspondingly greater than that for the smaller mass $M=10^{-18}$kg (four red lines). The figure also depicts scenarios at different temperatures, showing that the phonon-induced contrast loss at 300K is greater than that at 4K.
Note that since we start from a one-dimensional atomic chain model, the mass $M$ in (31) and (32) should, in principle, be understood as the mass of a single atomic chain within the diamond. However, in the subsequent calculations, we will interpret this $M$ as the total mass of an isotropic homogeneous cubic diamond with side length $L=(M/\rho)^{-\frac{1}{3}}$, where the main contribution to contrast loss comes from the fundamental tone $\omega_{0}=\pi c/L$.  [^4] Phonon modes depend on the shape of three-dimensional objects. Studies on acoustic phonons in circular objects can be found in Gonzalez-Ballestero et al. (2020). The situation becomes considerably more complex for arbitrary-shaped three-dimensional objects, and deserves separate discussion. .
@@
In Fig. 5, we show the contrast reduction resulting from spin-phonon (solid lines) and diamagnetism-phonon (dash and dot lines) interactions in a setup with $\Delta t=1$s. The horizontal axis represents the maximum separation distance of the diamond’s CoMs in the spatial superposition state. The black ($10^{-14}$ kg) and red ($10^{-18}$ kg) lines compare diamonds of different masses. We also compared the contrast loss at different temperatures, 555 Since the main contribution of contrast loss comes from the fundamental frequency $\omega_{0}\sim c/L$. One can consider two limits here, one is $\hbar\omega_{0}\ll k_{\text{B}}T$ and the other is $\hbar\omega_{0}\gg k_{\text{B}}T$. For the former limit, the square of the transfer function will yield the maximum value of $|\Gamma(\omega_{0})|^{2}\sim 1$, and Eqs. (31) and (32) become, respectively, $-\text{ln}~{}C_{\text{spin}}\approx\frac{2k_{\text{B}}T}{\hbar^{2}}\frac{M\Delta X^{2}_{\text{m}}}{\tau_{\text{a}}^{4}\omega_{0}^{4}}\sim\frac{2k_{\text{B}}T}{\hbar^{2}}\rho\left(\frac{\Delta X_{\text{m}}}{c^{2}\tau_{a}^{2}}\right)^{2}L^{7},$ $-\text{ln}~{}C_{\text{dia}}\approx\frac{k_{\text{B}}T\chi_{\rho}^{2}}{8\hbar^{2}\mu_{0}^{2}\mu^{4}}\frac{M^{5}\Delta X_{\text{m}}^{6}}{\tau_{\text{a}}^{12}\omega_{0}^{8}}\sim\frac{k_{\text{B}}T\chi_{\rho}^{2}\rho^{5}}{8\mu_{0}^{2}\mu^{4}\hbar^{2}}\left(\frac{\Delta X_{\text{m}}^{3}}{c^{4}\tau_{\text{a}}^{6}}\right)^{2}L^{23}.$ In the other limit, when $\hbar\omega_{0}\gg k_{\text{B}}T$ the hyperbolic function goes $\coth\frac{1}{2}\frac{\hbar\omega_{0}}{k_{\text{B}}T}\rightarrow 1.$ The contrast reductions from the two interactions become as, $-\text{ln}~{}C_{\text{spin}}\approx\frac{M\Delta X^{2}_{\text{m}}}{\hbar\tau_{\text{a}}^{4}\omega_{0}^{3}}\sim\frac{\rho}{\hbar c^{3}}\left(\frac{\Delta X_{\text{m}}L^{3}}{\tau_{\text{a}}^{2}}\right)^{2},$ $-\text{ln}~{}C_{\text{dia}}\approx\frac{\chi_{\rho}^{2}}{16\mu_{0}^{2}\mu^{4}\hbar}\frac{M^{5}\Delta X^{6}_{\text{m}}}{\tau_{\text{a}}^{12}\omega_{0}^{7}}\sim\frac{\chi_{\rho}^{2}}{16\mu_{0}^{2}\mu^{4}\hbar}\frac{\rho^{5}}{c^{7}}\left(\frac{\Delta X^{3}_{\text{m}}L^{11}}{\tau_{\text{a}}^{6}}\right)^{2}.$ In the above two limits, we have used the relation: $M=\rho L^{3}$, where $\rho$ is the density of test masses and $L$ is the length. In either of the limits, one can always conclude that more rigid material (which means faster sound speed) and smaller size of the test masses will produce less reduction in contrast. The diamond-like crystal will be an example of an ideal crystal for the QGEM experiment. where the results at low temperature (4K) are presented by dark-solid lines and dash lines while at high temperature (300K) results are shown in light-solid and dot lines. The results indicate that the contrast loss induced by the spin-phonon coupling is dominant, while the effect of diamagnetism-phonon interaction is significantly smaller. The source of these differences can be easily understood by comparisons in Eqs. (3132). The contrast reduction caused by the former interaction is suppressed by $\omega_{q}^{3}$, whereas the contrast loss caused by diamagnetism-phonon interaction is suppressed by $\omega_{q}^{7}$.
(a) (a)
(b) (b)
(c) (c)
(d) (d)
Figure 6: Panels (a), (b), (c), and (d) illustrate the contrast reduction resulting from spin-phonon coupling and diamagnetism-phonon interaction under different close-loop protocols, where the colour shading from dark to light represents progressively intensified contrast reduction. The first column corresponds to spin-phonon coupling, while the second corresponds to diamagnetism-phonon interaction. In the first row, (a) and (b) depict scenarios with a diamond mass of $M\sim 2.25\times 10^{-14}$ kg, showing different maximum separations $\Delta X_{\text{m}}$ (horizontal axis) and duration times $\Delta t$ (vertical axis). In the second row, (c) and (d) assume a desired separation distance of $\Delta X_{\text{m}}=10^{-4}$m for the diamond’s CoMs. We explore the contrast loss caused by phonon wave packets of different masses under varying duration times. In the four subfigures depicted, we have marked the cases where the contrast loss reaches $1\%$. In subfigures (a) and (b), for a diamond with mass $M=2.25\times 10^{-14}$kg, significant contrast loss requires achieving maximum separation distances of approximately $\sim 10^{-3}$m and $\sim 10^{-2}$m for the CoM wave packet within 0.05 s, respectively. In subfigures (c) and (d), to achieve $\Delta X_{\text{m}}=10^{-4}$m within 0.05 s, the diamond mass needs to reach approximately $\sim 10^{-13}$kg and $10^{-12}$kg, respectively, to exhibit significant contrast loss. In these four cases marked, the magnetic gradients required are all large, becoming approximately $1000$ to $10000$ times greater than the current experimental capability.
Moreover, room temperature does not drastically alter the outcomes for either of them. Additionally, it can be observed that larger diamond masses lead to more phonon contrast reduction, which is quite understandable. Larger mass implies a lower range of phonon frequencies. From Eqs. (26) and (29), it can be seen that these $\omega_{q}$ terms appear in the denominators.
@@
From Eq. (26), we can see that the acceleration protocol profoundly impacts the contrast of the phonon wave packets. In Fig. 6, we use a density plot to detail the contrast reduction resulting from different acceleration protocols. In the four subplots, we highlight the black curves, where the magnetic field gradient is precisely $10^{6}$ T/m, and in the region above it, the duration time is longer, thus requiring a smaller magnetic field gradient. In Fig. 6 (a) and (b), we study the contrast of a diamond with a mass of $M=2.25\times 10^{-14}$ kg under different maximum separation distances $\Delta X_{\text{m}}$ and duration times. In (c) and (d), we set the CoM’s maximum separation distance to $10^{-4}$ m which represent a sufficiently large separation distance of wave packets in QGEM and study the contrast loss for different diamond masses and duration times.
From (a) and (b), one can see that for $10^{-14}$kg diamond, when the separation protocol requires the loop to be completed within a short time (for example, 0.05 s) while reaching large separation distance $\Delta X_{\text{m}}$, then the loss of contrast from both interactions could go to $1\%$. Likewise, in (c) and (d), for large spatial separation as expected by the QGEM (for example $10^{-4}$m), protocols with heavier diamond and shorter duration time $\Delta t$ would also lead to severe contrast reduction in phonon degrees of freedom. Overall, significant contrast loss will happen when wave packets of a large mass diamond are separated over a large spatial distance within a short period. However, within the current experimental constraints where the SG magnetic field is constrained to be less than $10^{6}~{}{\rm T/m}$ Machluf et al. (2013); Mamin et al. (2012), the contrast loss induced by both couplings is negligible for QGEM.
@@ VII Discussion
In this paper, we have analyzed the contrast loss due to phonon in creating spatial superposition of a nanoscaled diamond embedded with a single NV centre based on a Stern-Gaelach interferometer. During the process of creating a spatial superposition state using the coupling between the spin and the linear magnetic field, the spin particle in the lattice chain generates a pair of forces (14) of opposite directions and equal magnitudes due to their own spin superposition state. Similar to the Humpty-Dumpty problem at the centre of mass of the diamond, the spin particles on both sides of the interferometer arm have different vibrational states in the lattice, resulting in differences in the quantum states of the phonons. This also leads to imperfections in the overlap of phonon wave packets.
The other interaction arises from the diamagnetic properties of diamond. Each atom in the lattice chain produces a diamagnetic force (15) when subjected to an external magnetic field. This diamagnetic force is of the same direction but has a different magnitude on both sides of the interferometer arm. Similarly, due to this difference, the quantum states of phonons on both sides are also different, resulting in contrast reduction.
In this paper, we use a linear magnetic field (9), which is widely used in atomic interferometers. Furthermore, this linearity greatly simplifies the calculation of the motion state of spin particles. That is, the splitting force is independent of trajectories and the equilibrium position.
The situation is more complex regarding the interaction between diamagnetic force and phonons. Since the diamagnetic term in the Hamiltonian includes $B(X_{i})^{2}$, the force on the lattice is now coupled with its trajectory. This external field mediated coupling between phonon and CoM’s trajectory is common in the study of optical cooling of the motion state of diamond centre. In this work, we typically studied the latter situation and made comparison of contrast reduction between these two interactions.
Specifically, we assumed that the phonons were initially at a specific temperature $T$, and we utilized the Wigner representation to express and compute the evolution of the quantum state of the phonons.
Our results indicate that among these two sources contributing to the contrast loss in the phonon degrees of freedom, the effect of the spin-phonon coupling is dominant. The contrast loss induced by the diamagnetic term is smaller than the former, as shown in Fig. 5. This result can be directly understood from (31) and (32). For the spin-phonon coupling, which does not depend on the trajectory, the contrast reduction is suppressed by the phonon frequency to the power of three. Meanwhile, the contrast loss due to diamagnetism-phonon interaction is suppressed by the phonon frequency to the power of seven.
The latter case directly arises from its coupling with the trajectory, as the diamagnetic force is a function of time, and its variation is adiabatic. In the seminal work Henkel and Folman (2022, 2023), this is also referred to as the degree of adiabaticity.
However, overall, the contrast reduction induced by both interactions is minimal. This is good news for matter-wave interferometers. There are two main reasons for this result. One is that we limit the gradient of the magnetic field. The most important reason, however, is that we are studying diamonds of very small size ($10^{-20}\sim 10^{-14}$ kg) compared to the speed of sound $c$ multiplied by the duration time (on the order of microseconds and 1 second for QGEM). Since the frequencies of the phonons are very high (tens of gigahertz to several terahertz), their contribution to contrast loss is tiny. However, if we prefer a different material where the speed of sound is much smaller than that of the diamond, we would see the loss of contrast occur. However, it also depends on how we create the superposition; hence, it depends on the details of the transfer function.
Acknowledgements: SB thanks EPSRC grants EP/R029075/1, EP/X009467/1, and ST/W006227/1. Q.X and R.Z are supported by the China Scholarship Council (CSC).
Appendix A Diamagnetic repulsion
From Eq. (8) one can immediately write down the diamagnetic repulsion given by the entire diamond,
$F_{\text{dia}}^{\text{L,R}}$ $=\partial_{X}\frac{\chi_{\rho}M}{2\mu_{0}}B^{2}(X)|_{X=X^{\text{L,R}}_{\text{c}}}$ (33)
$=\frac{\chi_{\rho}M}{\mu_{0}}(B_{0}\eta_{\text{b}}+\eta^{2}_{\text{b}}X^{\text{L,R}}_{\text{c}})$
The diamagnetic force here is a macroscopic quantity, representing the repulsion of the entire diamond to the magnetic field, which effectively located at CoM with equilibrium position $X_{\text{c}}$. However not like the splitting force which acts on the central particle, this does not mean that the repulsion exhibited by particles in the middle of the atomic chain is equal to Eq. (33).
Here, we consider that each atom in the atomic chain is magnetized with the potential energy $U(X_{i})$. We will not delve into the specific form of this $U(X_{i})$, but in general, it is the product of the magnetization of a single carbon atom and the magnetic field strength experienced by that atom (relevant discussions can be found in Chapter 4 of Kittel and McEuen (2018)). Therefore, each atom will contribute a different diamagnetic force, and these microscopic diamagnetic forces will vary in magnitude due to the different equilibrium positions of the atoms,
$F_{i,\text{dia}}^{\text{L,R}}$ $=-\partial_{X}U(X)|_{X=X^{\text{L,R}}_{i}}$ (34)
The summation over index $i$ of these microscopic diamagnetic forces must equal to Eq. (33), and thus we have the relation Eq. (15).
Appendix B Initial thermal equilibrium mixed state